# Derandomizing Multivariate Polynomial Factoring for Low Degree Factors

### v1

@@ 2.2 Transformation to a monic polynomial
$\tau_{{\boldsymbol{\alpha}}}:~{}~{}z_{i}\mapsto\alpha_{i}x+z_{i},$
@@
$nd\,\textup{T}_{{\rm PIT}(\textup{Hom}[\mathcal{C}])}+{\rm poly}(snd)\,.$
@@
Let $f({\boldsymbol{z}})\in\mathcal{C}$ be a polynomial of degree $d$ with $n$ variables ${\boldsymbol{z}}=(z_{1},\dots,z_{n})$
To see what the transformation does, let
$f=f_{0}+f_{1}+\cdots+f_{d},$
@@
$f_{d}({\boldsymbol{z}})=\sum_{|{\boldsymbol{\beta}}|_{1}=d}c_{{\boldsymbol{\beta}}}{\boldsymbol{z}}^{{\boldsymbol{\beta}}}.$
@@
Since we work with the shifted polynomial, we need to ensure that the shift of variables does not affect the irreducibility of the factors; this is guaranteed by the following lemma. It is quite standard in the literature; for a nice proof, see [KRSV24, Lemma B7].
@@ 2.3 Basics of factoring and interpolation and PIT
Berlekamp [Ber70] and Lenstra, Lenstra and Lovász [LLL82] gave efficient factorization algorithms for univariate polynomials over finite fields and $\mathbb{Q}$, respectively. Kaltofen [Kal85c] showed how to reduce the factorization of bivariate polynomials to univariate polynomials. In fact, the reduction works for $k$-variate polynomials, for any constant $k$. In our case, we use it for the case $k=3$.
@@
The following lemma shows how to find the multiplicity of an irreducible factor $g$ of a polynomial $f$. It holds when char$(\mathbb{F})=0$, or, large. For a concise proof, see [KRS24, Lemma 4.1].
@@
Let $f({\boldsymbol{z}}),g({\boldsymbol{z}})\in\mathbb{F}[{\boldsymbol{z}}]$ be non-zero polynomials and let $z\in\{z_{1},\cdots,z_{n}\}$ be such that $\partial_{z}(g)\neq 0$ and $g$ is irreducible. Then the multiplicity of $g$ in $f$ is the smallest non-negative integer $e$ such that $g\nmid\frac{\partial^{e}f}{\partial z^{e}}$.
Klivans and Spielman [KS01] derandomized the isolation lemma for PIT of sparse polynomials. Their algorithm works over fields of 00 or large characteristics.
Theorem 2.6 (Sparse PIT and interpolation [KS01]).
@@
$\textup{T}_{{\rm PIT}(\mathcal{C}_{\textup{sp}})}={\rm poly}(snd).$
@@
Limaye, Srinivasan, and Tavenas [LST21] designed a deterministic subexponential-time PIT for constant-depth circuits.
Theorem 2.7 (PIT for constant depth circuits [LST21, Corollary 6]).
@@
$\textup{T}_{{\rm PIT}(\mathcal{C}_{\textup{Depth-}{t}})}=\left(ns^{O(t)}\right)^{O((sd)^{\varepsilon})}.$
@@ 2.4 Divisibility testing reduces to PIT
When $f,g$ can be computed by circuits of size $s$, Strassen [Str73] showed that if $g|f$, then $h=f/g$ can be computed by a circuit of size ${\rm poly}(sd)$, where $d=\deg(h)$. Forbes [For15] observed that even in the case when $g\not|\,f$, one can follow Strassen’s argument and obtain a small size circuit that computes a polynomial $\widetilde{h}$ such that $g|f\iff f=g\widetilde{h}$. Hence, we have a reduction from divisibility testing to PIT.
Lemma 2.8 (Divisibility reduces to PIT [For15, Corollary 7.10]).
@@
$g({\boldsymbol{z}})|f({\boldsymbol{z}})\iff f({\boldsymbol{z}}+{\boldsymbol{\alpha}})=g({\boldsymbol{z}}+{\boldsymbol{\alpha}})\,\widetilde{h}({\boldsymbol{z}}).$
@@
$\textup{T}_{{\rm Div}(\mathcal{C}_{\textup{Depth-}{t}}/{\mathcal{D}}_{\textup{sp}})}=\left(n\,(sd)^{O(t)}\right)^{O((sd)^{\epsilon})}\,.$
@@
$\widetilde{f}=f({\boldsymbol{z}}+{\boldsymbol{\alpha}})-g({\boldsymbol{z}}+{\boldsymbol{\alpha}})\cdot\widetilde{h}({\boldsymbol{z}})$
@@
We also consider divisibility of sparse polynomials by constant-degree polynomials. Building on Lemma 2.8, Forbes [For15] reduced the problem to a PIT that can be solved in quasipolynomial time.
Corollary 2.10 (Sparse by constant degree division [For15, Corollary 7.17]).
@@
$\textup{T}_{{\rm Div}(\mathcal{C}_{\textup{sp}}/{\mathcal{D}}_{\delta})}=(snd)^{O(\delta\log s)}\,.$
@@ 2.5 Effective Hilbert’s Irreducibility Theorem
Factorization algorithms often start with an effective version of Hilbert’s Irreducibility Theorem due to Kaltofen and von zur Gathen. It shows how to project a multivariate irreducible polynomial down to two variables, such that the projected bivariate polynomial stays irreducible. The proof shows the existence of an irreducibility certifying polynomial $G({\boldsymbol{b}},{\boldsymbol{c}})$ in $2n$ variables corresponding to the irreducible polynomial $g(x,{\boldsymbol{z}})$. The nonzeroness of $G$ proves the irreducibility of $g(x,{\boldsymbol{z}})$ and also gives a way to find an irreducibility-preserving projection to bivariate (see [Kal85b, Kal95, KSS15]).
@@
$G({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\neq 0\implies\widehat{g}(x,t)=g(x,{\boldsymbol{\beta}}t+{\boldsymbol{\gamma}})\text{ is irreducible,}$, (10)
@@
We define a corresponding computational problem. Let $\mathcal{C}\subseteq{\mathcal{P}}(n,d)$ be a class of polynomials and $g({\boldsymbol{z}})\in\mathcal{C}$. Assume we have already computed an ${\boldsymbol{\alpha}}\in(\mathbb{F}\setminus\{0\})$ as in Lemma 2.2 that the shifted polynomial $g_{{\boldsymbol{\alpha}}}(x,{\boldsymbol{z}})$ is monic. Now we want to find a hitting set according to (10).
@@
    Given ${\boldsymbol{\alpha}}\in(\mathbb{F}\setminus\{0\})^{n}$, compute a set $H_{{\boldsymbol{\alpha}}}\subseteq\mathbb{F}^{2n}$ such that for all $g\in\mathcal{C}$ where $g({\boldsymbol{\alpha}}x+{\boldsymbol{z}})$ is monic in $x$, we have
    $g\text{ irreducible }\implies\exists({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in H_{{\boldsymbol{\alpha}}}~{}~{}~{}g({\boldsymbol{\alpha}}x+{\boldsymbol{\beta}}t+{\boldsymbol{\gamma}})\in\mathbb{F}[x,t]~{}\text{ is irreducible}.$
@@ 2.6 Isolation
$M_{\delta}=\left\{\,{\boldsymbol{z}}^{{\boldsymbol{e}}}\mid||{\boldsymbol{e}}||_{1}\leq\delta\,\right\}.$ (11)
@@
$|M_{\delta}|\leq\binom{n+\delta}{\delta}\leq(n+\delta)^{\delta}\leq(\delta+1)\,n^{\delta}=O(n^{\delta}).$ (12)
@@
Clearly, $\varphi$ is injective on $M_{\delta}$. However, the degree of $y$ can be exponentially large, up to $(\delta+1)^{n}$. A way around is to take the exponents modulo some small prime number $p$. We have to determine $p$ in a way to keep the mapping injective on $M_{\delta}$. Hence, for any two terms $y^{e},y^{e^{\prime}}$ we get from $\varphi$, we have to ensure that $e\not\equiv e^{\prime}\pmod{p}$. Equivalently $p\not|\,(e-e^{\prime})$.
We have $|e-e^{\prime}|\leq(\delta+1)^{n}$ and, by (12), there are $(\delta+1)^{2}n^{2\delta}$ many pairs $e,e^{\prime}$ we get from $M_{\delta}$ via $\varphi$. Prime $p$ should not divide any of these differences, and hence, $p$ should not divide their product $P$. The product $P$ is bounded by
$P\leq\left((\delta+1)^{n}\right)^{(\delta+1)^{2}n^{2\delta}}=(\delta+1)^{(\delta+1)^{2}n^{2\delta+1}}.$ (15)
Hence, $P$ has at most $\log P\leq R=(\delta+1)^{3}\,n^{2\delta+1}$ many prime factors. By the Prime Number Theorem, there are more than $\log P$ primes in the set $[R^{2}]$. Hence, we can find an appropriate prime $p\leq R^{2}=n^{O(\delta)}$.
@@
Note that in Lemma 2.13, we do not claim that irreducibility is maintained: when $g$ is irreducible, still $\varphi_{p}(g)$ might be reducible. Consider the example $n=\delta=2$. The weights $\{1,3\}$ make sure that each monomial $z_{1}^{2},z_{1}z_{2},z_{2}^{2}$ gets mapped to a distinct power in $y$. Let $g(x,{\boldsymbol{z}})=x^{2}-z_{1}z_{2}$. Observe that $g$ is irreducible, however $g(x,y,y^{3})=(x-y^{2})(x+y^{2})$ is reducible.
@@
$\widehat{G}(y)=G(y^{w_{1}},\dots,y^{w_{n}},y^{w^{\prime}_{1}},\dots,y^{w^{\prime}_{n}})\not=0\;.$
@@
To compute $g$ from $\Psi(g)$, set $t=0$, i.e. consider $\Psi(g)(x,y,0)$. From (18) we see that monomials then have the form
$c_{k,{\boldsymbol{e}}}\,x^{k}\,y^{\sum_{i=1}^{n}e_{i}w^{\prime}_{i}}\,.$
From these we get the exponents $k$ and ${\boldsymbol{e}}$ similar as in the proof of Lemma 2.12. ∎
Remark.
In Corollary 2.14, when we say that we invert $\Psi$, it means that for a given $h\in\mathbb{F}[x,y,t]$ which is monic in $x$ with $x$-degree $\leq\delta$, we either detect that $h$ is not in the codomain of $\Psi$, or we compute $g\in\mathbb{F}[x,{\boldsymbol{z}}]$ such that $\Psi(g)=h$ in time $n^{{\rm poly}(\delta)}$.
The inversion can be done similarly as described in the proof of Corollary 2.14. One can evaluate $t=0$, and then for every monomial $x^{k}y^{j}$, try to find $x^{k}{\boldsymbol{z}}^{{\boldsymbol{e}}}$ that would map to such a monomial at $t=0$. By the property of the map, while mapping the $y$-degrees, ${\boldsymbol{z}}$-degree could be at most $\delta$, i.e. $\deg(x^{k}{\boldsymbol{z}}^{{\boldsymbol{e}}})\leq 2\delta$. We will, of course, return empty if the degree of any such monomial, after inverting, becomes $>\delta$. Finally, once we have got a candidate $g$ of degree $\delta$, we still have to check whether $\Psi(g)=h$, because the inversion procedure ignores the variable $t$. The last step can also be efficiently checked.
@@
Finally, we conclude this subsection by a general remark that whenever $n$ and $\delta$ are fixed, these weights are fixed and can be found efficiently.
@@ 3.1 Projected constant degree factors
The next step is to factor $3$-variate $\Psi(f_{{\boldsymbol{\alpha}}})$ via Lemma 2.4. The following pseudo-code summarizes the steps.
@@
Output : All projected factors of $f$ of degree $\delta$ with their multiplicities.
@@
3
4Find ${\boldsymbol{w}},{\boldsymbol{w}}^{\prime}\in\mathbb{F}^{n}$ as in Corollary 2.15 and compute $\Psi(f_{{\boldsymbol{\alpha}}})\in\mathbb{F}[x,y]$ in dense representation
@@
5
6Define $S_{\textup{Proj-Fac}}=\left\{\,h_{i}\mid\deg_{x}(h_{i})\leq\delta\,\right\}$ and $S_{\textup{Proj-Fac-mult}}=\left\{\,(h_{i},e_{i})\mid h_{i}\in S_{\textup{Proj-Fac}}\,\right\}$

### v2

@@ 2.2 Transformation to a monic polynomial
$\tau_{{\boldsymbol{\alpha}}}:~~z_{i}\mapsto\alpha_{i}x+z_{i},$ (11)
@@
$nd\,\textup{T}_{{\rm PIT}(\textup{Hom}[\mathcal{C}])}+{\rm poly}(snd)\,.$ (12)
@@
Let $f({\boldsymbol{z}})\in\mathcal{C}$ be a polynomial of degree $d$ with $n$ variables ${\boldsymbol{z}}=(z_{1},z_{2},\dots,z_{n})$. To see what the transformation does, let
$f=f_{0}+f_{1}+\cdots+f_{d},$ (13)
@@
$f_{d}({\boldsymbol{z}})=\sum_{|{\boldsymbol{\beta}}|_{1}=d}c_{{\boldsymbol{\beta}}}{\boldsymbol{z}}^{{\boldsymbol{\beta}}}.$ (14)
@@
Since we work with the shifted polynomial, we need to ensure that the shift of variables does not affect the irreducibility of the factors; this is guaranteed by the following lemma. It is quite standard in the literature; for a nice proof, see [kumar2024towards, Lemma B7].
@@ 2.3 Basics of factoring and interpolation and PIT
Berlekamp [berlekamp1970factoring] and Lenstra, Lenstra and Lovász [lenstra1982factoring] gave efficient deterministic factorization algorithms for univariate polynomials over finite fields and $\mathbb{Q}$, respectively. Kaltofen [kaltofen1985polynomial] showed how to reduce the factorization of bivariate polynomials to univariate polynomials. In fact, the reduction works for $k$-variate polynomials, for any constant $k$. In our case, we use it for the case $k=3$.
@@
[…]
@@
Let $f({\boldsymbol{z}}),g({\boldsymbol{z}})\in\mathbb{F}[{\boldsymbol{z}}]$ be non-zero polynomials and let $z\in\{z_{1},z_{2},\cdots,z_{n}\}$ be such that $\partial_{z}(g)\neq 0$ and $g$ is irreducible. Then the multiplicity of $g$ in $f$ is the smallest non-negative integer $e$ such that $g\nmid\frac{\partial^{e}f}{\partial z^{e}}$.
Klivans and Spielman [klivans2001randomness] derandomized the isolation lemma for PIT of sparse polynomials. Their algorithm works over fields of 0 or large characteristic.
Theorem 2.6 (Sparse PIT and interpolation [klivans2001randomness]).
@@
$\textup{T}_{{\rm PIT}(\mathcal{C}_{\textup{sp}})}={\rm poly}(snd).$ (15)
@@
Limaye, Srinivasan, and Tavenas [limaye2022superpolynomial] designed a deterministic subexponential-time PIT for constant-depth circuits.
Theorem 2.7 (PIT for constant depth circuits [limaye2022superpolynomial, Corollary 6]).
@@
$\textup{T}_{{\rm PIT}(\mathcal{C}_{\textup{Depth-}{t}})}=\left(ns^{O(t)}\right)^{O((sd)^{\varepsilon})}.$ (16)
@@ 2.4 Divisibility testing reduces to PIT
[…]
Lemma 2.8 (Divisibility reduces to PIT [forbes2015deterministic, Corollary 7.10]).
@@
$g({\boldsymbol{z}})|f({\boldsymbol{z}})\iff f({\boldsymbol{z}}+{\boldsymbol{\alpha}})=g({\boldsymbol{z}}+{\boldsymbol{\alpha}})\,\widetilde{h}({\boldsymbol{z}}).$ (18)
@@
$\textup{T}_{{\rm Div}(\mathcal{C}_{\textup{Depth-}{t}}/{\mathcal{D}}_{\textup{sp}})}=\left(n\,(sd)^{O(t)}\right)^{O((sd)^{\epsilon})}\,.$ (19)
@@
$\widetilde{f}=f({\boldsymbol{z}}+{\boldsymbol{\alpha}})-g({\boldsymbol{z}}+{\boldsymbol{\alpha}})\cdot\widetilde{h}({\boldsymbol{z}})$ (20)
@@
We also consider divisibility of sparse polynomials by constant-degree polynomials. Building on Lemma 2.8, Forbes [forbes2015deterministic] reduced the problem to a PIT that can be solved in quasi-polynomial time.
Corollary 2.10 (Sparse by constant degree division [forbes2015deterministic, Corollary 7.17]).
@@
$\textup{T}_{{\rm Div}(\mathcal{C}_{\textup{sp}}/{\mathcal{D}}_{\delta})}=(snd)^{O(\delta\log s)}\,.$ (21)
@@ 2.5 Effective Hilbert’s Irreducibility Theorem
Factorization algorithms often start with an effective version of Hilbert’s Irreducibility Theorem due to Kaltofen and von zur Gathen. It shows how to project a multivariate irreducible polynomial down to two variables, such that the projected bivariate polynomial stays irreducible. The proof shows the existence of an irreducibility certifying polynomial $G({\boldsymbol{b}},{\boldsymbol{c}})$ in $2n$ variables corresponding to the irreducible polynomial $g(x,{\boldsymbol{z}})$. The nonzeroness of $G$ proves the irreducibility of $g(x,{\boldsymbol{z}})$ and also gives a way to find an irreducibility-preserving projection to bivariate (see [kaltofen1985effective, kaltofen1995effective, kopparty2015equivalence]).
@@
$G({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\neq 0\implies\widehat{g}(x,t)=g(x,{\boldsymbol{\beta}}t+{\boldsymbol{\gamma}})\text{ is irreducible,}$ (22)
@@
We define a corresponding computational problem. Let $\mathcal{C}\subseteq{\mathcal{P}}(n,d)$ be a class of polynomials and $g({\boldsymbol{z}})\in\mathcal{C}$. Assume we have already computed an ${\boldsymbol{\alpha}}\in(\mathbb{F}\setminus\{0\})$ as in Lemma 2.2 that the shifted polynomial $g_{{\boldsymbol{\alpha}}}(x,{\boldsymbol{z}})$ is monic. Now we want to find a hitting set for $g_{{\boldsymbol{\alpha}}}$ according to (22).
@@
    Given ${\boldsymbol{\alpha}}\in(\mathbb{F}\setminus\{0\})^{n}$, compute a set $H_{{\boldsymbol{\alpha}}}\subseteq\mathbb{F}^{2n}$ such that for all $g\in\mathcal{C}$, where $g({\boldsymbol{\alpha}}x+{\boldsymbol{z}})$ is monic in $x$, we have
    $g\text{ irreducible }\implies\exists({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in H_{{\boldsymbol{\alpha}}}~~~g({\boldsymbol{\alpha}}x+{\boldsymbol{\beta}}t+{\boldsymbol{\gamma}})\in\mathbb{F}[x,t]~\text{ is irreducible}.$ (23)
@@
Remark.
A subtlety in the definition of ${\rm Irred\mbox{-}Proj}(\mathcal{C})$ is that it asks to compute a hitting set $H_{{\boldsymbol{\alpha}}}$ for the shifted polynomial $g_{{\boldsymbol{\alpha}}}$. If $\mathcal{C}$ is closed under variable shifts, then the hitting set is for polynomials in $\mathcal{C}$. However, for example $\mathcal{C}_{\textup{sp}}$ is not closed under variable shifts. Hence, ${\rm Irred\mbox{-}Proj}(\mathcal{C}_{\textup{sp}})$ asks to compute a hitting set for a non-sparse polynomial that comes from a shift of a sparse polynomial.
@@ 2.6 Isolation
[…]
@@
$|M_{\delta}|\leq\binom{n+\delta}{\delta}\leq(n+\delta)^{\delta}\leq(\delta+1)\,\delta^{\delta}\,n^{\delta}=O(n^{\delta}).$ (25)
@@
[…]
We have $|e-e^{\prime}|\leq(\delta+1)^{n}$ and, by (25), there are $O(n^{2\delta})$ many pairs $e,e^{\prime}$ we get from $M_{\delta}$ via $\varphi$. Prime $p$ should not divide any of these differences, and hence, $p$ should not divide the product $P$ of all these differences. The product $P$ is bounded by
$P\leq\left((\delta+1)^{n}\right)^{O(n^{2\delta})}=(\delta+1)^{O(n^{2\delta+1})}.$ (28)
Hence, $P$ has at most $\log P\leq R=O(n^{2\delta+1})$ many prime factors. By the Prime Number Theorem, there are more than $\log P$ primes in $R^{2}]$. Hence, we can find an appropriate prime $p\leq R^{2}=n^{O(\delta)}$.
@@
Note that in Lemma 2.13, we do not claim that irreducibility is maintained: when $g$ is irreducible, still $\varphi_{p}(g)$ might be reducible. Consider the example $n=\delta=2$. The weights $\{1,3\}$ make sure that each monomial in $M_{2}=\{\,z_{1},~z_{2},~z_{1}^{2},~z_{1}z_{2},~z_{2}^{2}\,\}$ gets mapped to a distinct power in $y$. Let $g({\boldsymbol{z}})=1-z_{1}z_{2}$. Observe that $g$ is irreducible, however $g(y,y^{3})=(1-y^{2})(1+y^{2})$ is reducible.
@@
$\widehat{G}(y)=G(y^{w_{1}},\dots,y^{w_{n}},y^{w^{\prime}_{1}},\dots,y^{w^{\prime}_{n}})\not=0\;.$ (31)
@@
For inversion, we have given $h\in\mathbb{F}[x,y,t]$ which is monic in $x$ with $x$-degree $\leq\delta$. We either have to compute $g$ such that $\Psi(g)=h$, or detect that $h$ is not in the codomain of $\Psi$.
We set $t=0$, i.e. we consider $h(x,y,0)$. From (32) we see that monomials then must have the form
$c_{k,{\boldsymbol{e}}}\,x^{k}\,y^{\sum_{i=1}^{n}e_{i}w^{\prime}_{i}}\,.$ (33)
From these we get the exponents $k$ and $\boldsymbol{e}$} for monomial $x^{k}\,{\boldsymbol{z}}^{{\boldsymbol{e}}}$, similar as in the proof of Lemma 2.12. In case we thereby get a degree $\delta$, then $h$ is not in the codomain of $\Psi$. Otherwise, we have got a candidate $g$ of degree $\delta$. But we still have to check whether $\Psi(g)=h$, because the inversion procedure ignores the variable $t$. The running time for inversion is $n^{{\rm poly}(\delta)}$. ∎
@@ 3.1 Projected constant degree factors
The next step is to factor $3$-variate $\Psi(f_{{\boldsymbol{\alpha}}})$ via Lemma 2.4. Algorithm 1, Projected-Factoring, below summarizes the steps.
@@
Output : All projected 33-variate factors of $f$ of degree $\delta$ with their multiplicities.
@@
4
5Find ${\boldsymbol{w}},{\boldsymbol{w}}^{\prime}\in\mathbb{F}^{n}$ as in Corollary 2.15 and compute $\Psi(f_{{\boldsymbol{\alpha}}})\in\mathbb{F}[x,y]$ in dense representation
@@
6
7Define $S_{\textup{Proj-Fac}}=\left\{\,h_{i}\mid\deg_{x}(h_{i})\leq\delta\,\right\}$ and $S_{\textup{Proj-Fac-mult}}=\left\{\,(h_{i},e_{i})\mid h_{i}\in S_{\textup{Proj-Fac}}\,\right\}$